OpenStudy (majorwhipage1):

The equation of line CD is y =-2x -2 write an equation,of a line parallel to line cd in slope-intercept form that contains (4,5) Y=-2x +13 Y=-1/2x+7 Y=-1/2x+3 Y=-2x-3

OpenStudy (queenie):

a parallel line will have the same slope...so can u tell me the slope of the given equation ?

OpenStudy (majorwhipage1):

-2x-2

OpenStudy (queenie):

not quite. In y = mx + b, the slope is in the m position. y = mx + b y = -2x - 2 as u can see, the slope is -2 understand so far ?

OpenStudy (majorwhipage1):

Yes

OpenStudy (queenie):

now we use y = mx + b slope(m) = -2 (4,5)...x = 4 and y = 5 now we sub...we are looking for b, the y intercept y = mx + b 5 = -2(4) + b can u finish and solve for b ?

OpenStudy (majorwhipage1):

5=-8+b 13=b Answer is y=-2x +13

OpenStudy (queenie):

correct

OpenStudy (majorwhipage1):

Thank you so much

OpenStudy (queenie):

yw